### v2

@@ Abstract
We propose a randomized lattice algorithm for approximating multivariate periodic functions over the $d$-dimensional unit cube from the weighted Korobov space with mixed smoothness $\alpha>1/2$ and product weights $\gamma_{1},\gamma_{2},\ldots\in[0,1]$. Building upon the deterministic lattice algorithm by Kuo, Sloan, and Woźniakowski (2006), we incorporate a randomized quadrature rule by Dick, Goda, and Suzuki (2022) to accelerate the convergence rate. This randomization involves drawing the number of points for function evaluations randomly, and selecting a good generating vector for rank-1 lattice points using the randomized component-by-component algorithm. We prove that our randomized algorithm achieves a worst-case root mean squared $L_{2}$-approximation error of order $M^{-\alpha/2-1/8+\varepsilon}$ for an arbitrarily small $\varepsilon>0$, where $M$ denotes the maximum number of function evaluations, and that the error bound is independent of the dimension $d$ if the weights satisfy $\sum_{j=1}^{\infty}\gamma_{j}^{1/\alpha}<\infty$. Our upper bound converges faster than a lower bound on the worst-case $L_{2}$-approximation error for deterministic rank-1 lattice-based approximation proved by Byrenheid, Kämmerer, Ullrich, and Volkmer (2017). We also show a lower error bound of order $M^{-\alpha/2-1/2}$ for our randomized algorithm, leaving a slight gap between the upper and lower bounds open for future research.
@@ 1 Introduction
where $\|\cdot\|_{d,\alpha,\boldsymbol{\gamma}}$ denotes the norm of the weighted Korobov space $H_{d,\alpha,\boldsymbol{\gamma}}$ and $\|\cdot\|_{L_{2}}$ denotes the $L_{2}$ norm over $[0,1)^{d}$.
@@
Following [7], in our randomized lattice-based algorithm for multivariate $L_{2}$-approximation, we draw the number of points $N$ randomly from a set of primes in the interval $(\lceil M/2\rceil,M]$ for a given $M$ and then select a generating vector using a randomized CBC algorithm. We prove that our randomized lattice-based algorithm attains the worst-case RMSE of order $M^{-\alpha/2-1/8+\varepsilon}$, and that the error bound is independent of the dimension $d$ if the weight parameters satisfy $\sum_{j=1}^{\infty}\gamma_{j}^{1/\alpha}<\infty$. It is striking that the convergence rate we obtain is better than the lower bound for any deterministic lattice-based algorithm as proven by Byrenheid, Kämmerer, Ullrich, and Volkmer [3]. Furthermore, we prove a lower error bound of order $M^{-\alpha/2-1/2}$ for our randomized algorithm. Filling the gap between the upper and lower bounds is left open for future research.
@@ 2.1 Lattice rules
[…]
@@ 2.3 Lattice algorithm for approximation
As a reference value, we use the initial error where we approximate the function with 00, i.e.,
@@
see [8, Section 3.4] and [9, Remark 13.4]. When $\alpha$ is an integer, we further have
@@ 3 Construction algorithm
In the step 3.(b), we need to arrange the integers $11,\ldots,N-$ 1 such that the corresponding value $R_{N,s,\alpha,\boldsymbol{\gamma}}$ is listed in ascending order and then pick one of the first $\lceil\tau(N-1)\rceil$ integers. This arrangement may not be unique if some of the integers yield the same value of $R_{N,s,\alpha,\boldsymbol{\gamma}}$. However, we can always make the ordering unique by further arranging those integers in ascending order. Moreover, as shown in Remark 2.5, when $\alpha$ is an integer, $R_{N,s,\alpha,\boldsymbol{\gamma}}$ has a computable formula as
@@
For the induction step, let $\boldsymbol{z}_{s-1}$ be the $(s-1)$-dimensional vector drawn by the first $s-1$ steps of the algorithm. For any $z_{s}\in\{1,\ldots,N-1\}$, by separating the cases with $\ell_{s}=0$ and $\ell_{s}\neq 0$, we have
@@
×$\quad\quad\quad\times\sum_{\boldsymbol{h}\in\mathbb{Z}^{s-1}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{s-1}\\ \boldsymbol{\ell}\cdot\boldsymbol{z}_{s-1}\equiv-\ell_{s}z_{s}\ (\mathrm{mod}\ N)\end{subarray}}\frac{1}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}$
@@
×$\quad\quad\quad\times\sum_{\boldsymbol{h}\in\mathbb{Z}^{s-1}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{s-1}\\ \boldsymbol{\ell}\cdot\boldsymbol{z}_{s-1}\equiv-\ell_{s}z_{s}\ (\mathrm{mod}\ N)\end{subarray}}\frac{1}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}.$
@@
×$\quad\quad\quad\times\sum_{\boldsymbol{h}\in\mathbb{Z}^{s-1}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{s-1}\\ \boldsymbol{\ell}\cdot\boldsymbol{z}_{s-1}\equiv-\ell_{s}z_{s}\ (\mathrm{mod}\ N)\end{subarray}}\frac{1}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}))^{1/\lambda}}$
@@
×$\quad\quad\quad\times\sum_{\boldsymbol{h}\in\mathbb{Z}^{s-1}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{s-1}\\ \boldsymbol{\ell}\cdot\boldsymbol{z}_{s-1}\equiv 0\ (\mathrm{mod}\ N)\end{subarray}}\frac{1}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}))^{1/\lambda}}$
@@
×$\quad\quad\quad\times\sum_{\boldsymbol{h}\in\mathbb{Z}^{s-1}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{s-1}\\ \boldsymbol{\ell}\cdot\boldsymbol{z}_{s-1}\not\equiv 0\ (\mathrm{mod}\ N)\end{subarray}}\frac{1}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}))^{1/\lambda}}$
@@ 4.1 Our algorithm
where $N,\boldsymbol{z},\boldsymbol{\Delta}$ are the random variables and $\boldsymbol{\Delta}$ follows the uniform distribution over $[0,1)^{d}$, and
@@
$e^{\mathrm{rms}\text{-}L_{2}\text{-}\mathrm{app}}_{d,\alpha,\boldsymbol{\gamma}}(A^{\mathrm{rand}}_{N,\boldsymbol{z},\boldsymbol{\Delta},\mathcal{A}_{d}(T)})$
@@
Although our randomized method is based on Algorithm 3.1, it can be substituted with Algorithm 2.4 in [7]. The difference lies in the error criterion being minimized. While we omit detailed discussion in this paper, the resulting randomized method based on Algorithm 2.4 in [7] leads to a weaker bound on the worst-case RMSE. Specifically, the method achieves a rate arbitrarily close to $M^{-(\alpha+1/2)/(2+1/(2\alpha))}$ for $\alpha>1/2$. This rate is inferior to the one we show in the next subsection, which is arbitrarily close to $M^{-\alpha/2-1/8}$. This difference in convergence rates, arising from the choice of error criterion, is analogous to the difference in convergence rates observed between the two deterministic lattice-based methods in [9, Chapter 13]—namely, “A direct approach” and “Improving the convergence rate”—which are likewise based on different quality criteria.
@@ 4.2 Error Analysis
where $\mathbb{I}_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}}^{\perp}}$ denotes the indicator function that returns $1$ if $\boldsymbol{\ell}\in P_{N,\boldsymbol{z}}^{\perp}$ holds and 00 otherwise. Note that $\omega(\boldsymbol{\ell})$ is nothing but the probability that a given frequency $\boldsymbol{\ell}$ belongs to the dual lattice $P_{N,\boldsymbol{z}}^{\perp}$ when $\boldsymbol{z}$ follows the uniform distribution over $\mathcal{Z}_{N,d,\tau}$ given $N$, and $N$ follows the uniform distribution over $\mathcal{P}_{M}$. Note that the size of $\mathcal{Z}_{N,d,\tau}$ can be bounded below as
@@
$H_{M}:=\inf_{1/2<\lambda<\alpha}\left(\frac{2}{(1-\tau)M}\prod_{j=1}^{d}\left(1+2^{2\alpha+2}\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{2}\right)^{\lambda}.$ (8)
@@
$M\geq\inf_{1/2<\lambda<\alpha}\frac{2}{1-\tau}\prod_{j=1}^{d}\left(1+2^{2\alpha+2}\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{2}$ (9)
@@
    There exists a constant $c>$ such that, for all $\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}$ we have
@@
$\leq\frac{1}{T}+T\frac{C_{\alpha,\beta,\lambda,\tau}}{M^{2\lambda-\lambda\beta/\alpha+1/2}}\prod_{j=1}^{d}\left(1+2^{2\alpha+2}\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{4\lambda-2\lambda\beta/\alpha}\max(1,2^{\beta}\gamma_{j}^{\beta/\alpha}),$
for any $\lambda\in(1/2,\alpha)$ and $\beta\in(0,\min(\alpha(1-1/(2\lambda)),1))$, with a constant $C_{\alpha,\beta,\lambda,\tau}>0$ independent of $M,d,\boldsymbol{\gamma},T$.
@@
$\left(e^{\mathrm{rms}\text{-}L_{2}\text{-}\mathrm{app}}_{d,\alpha,\boldsymbol{\gamma}}(A^{\mathrm{rand}}_{N,\boldsymbol{z},\boldsymbol{\Delta},\mathcal{A}_{d}(T)})\right)^{2}$ $\leq\frac{1}{T}+\sum_{\boldsymbol{h}\in\mathcal{A}_{d}(T)}\frac{T}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})}\sup_{\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}\frac{\omega(\boldsymbol{\ell})}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}$
$\leq\frac{1}{T}+T\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}\sup_{\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}\frac{\omega(\boldsymbol{\ell})}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}.$
@@
Using this and the first assertion of Lemma 4.4, together with the elementary inequality $\log(1+x)\leq x^{\beta}/\beta$ for $x>0$ and $10<\beta<$ 1 with $\log$ denoting the natural logarithm, we have
$B_{M}^{2}$ $=\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}\sup_{\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}\frac{\omega(\boldsymbol{\ell})}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}$
$\leq\frac{c}{\tau M}\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}\sup_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\geq H^{-1}_{M}\end{subarray}}\frac{\log\left(1+\|\boldsymbol{\ell}\|_{\infty}\right)}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}$
$\leq\frac{c}{\beta\tau M}\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}\sup_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\geq H^{-1}_{M}\end{subarray}}\frac{\|\boldsymbol{\ell}\|_{\infty}^{\beta}}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}$
$\leq\frac{c}{\beta\tau M}\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}\sup_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\geq H^{-1}_{M}\end{subarray}}\frac{\left(r_{\alpha,\boldsymbol{\gamma}}^{2}(\boldsymbol{\ell})\right)^{\beta/(2\alpha)}}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})},$
@@
$B_{M}^{2}$ $\leq\frac{c}{\beta\tau M}\prod_{j=1}^{d}\max(1,2^{\beta}\gamma_{j}^{\beta/\alpha})\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}\sup_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\geq H^{-1}_{M}\end{subarray}}\frac{1}{\left(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\right)^{2-\beta/\alpha}}.$
To simplify the notation, let us write
$\theta_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}):=\inf_{\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}),$
and, for $m\in\mathbb{N},$
$A_{\alpha,\boldsymbol{\gamma}}(m):=\sum_{\begin{subarray}{c}\boldsymbol{h}\in\mathbb{Z}^{d}\\ \theta_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})<m\end{subarray}}1.$
Then, for the sum over $\boldsymbol{h}$ above, we obtain

### v3

@@ Abstract
We propose a randomized lattice algorithm for approximating multivariate periodic functions over the $d$-dimensional unit cube from the weighted Korobov space with mixed smoothness $\alpha>1/2$ and product weights $\gamma_{1},\gamma_{2},\ldots\in[0,1]$. Building upon the deterministic lattice algorithm by Kuo, Sloan, and Woźniakowski (2006), we incorporate a randomized quadrature rule by Dick, Goda, and Suzuki (2022) to accelerate the convergence rate. This randomization involves drawing the number of points for function evaluations randomly, and selecting a good generating vector for rank-1 lattice points using the randomized component-by-component algorithm. We prove that our randomized algorithm achieves a worst-case root mean squared $L_{2}$-approximation error of order $M^{-\alpha(2\alpha+1)/(4\alpha+1)+\varepsilon}$ for an arbitrarily small $\varepsilon>0$, where $M$ denotes the maximum number of function evaluations, and that the error bound is independent of the dimension $d$ if the weights satisfy $\sum_{j=1}^{\infty}\gamma_{j}^{1/\alpha}<\infty$. Our upper bound converges faster than a lower bound on the worst-case $L_{2}$-approximation error for deterministic rank-1 lattice-based approximation proved by Byrenheid, Kämmerer, Ullrich, and Volkmer (2017). We also show a lower error bound of order $M^{-\alpha/2-1/2}$ for our randomized algorithm, leaving a slight gap between the upper and lower bounds open for future research.
@@ 1 Introduction
where $\|\cdot\|_{d,\alpha,\boldsymbol{\gamma}}$ denotes the norm of the weighted Korobov space $H_{d,\alpha,\boldsymbol{\gamma}}$ and $\|\cdot\|_{L_{2}}$ denotes the $L_{2}$ norm over $0,1)^{d}$.
@@
Following [7], in our randomized lattice-based algorithm for multivariate $L_{2}$-approximation, we draw the number of points $N$ randomly from a set of primes in the interval $(\lceil M/2\rceil,M]$ for a given $M$ and then select a generating vector using a randomized CBC algorithm. We prove that our randomized lattice-based algorithm attains the worst-case RMSE of order $M^{-\alpha(2\alpha+1)/(4\alpha+1)+\varepsilon}$, and that the error bound is independent of the dimension $d$ if the weight parameters satisfy $\sum_{j=1}^{\infty}\gamma_{j}^{1/\alpha}<\infty$. It is striking that the convergence rate we obtain is better than the lower bound for any deterministic lattice-based algorithm as proven by Byrenheid, Kämmerer, Ullrich, and Volkmer [3]. Furthermore, we prove a lower error bound of order $M^{-\alpha/2-1/2}$ for our randomized algorithm. Filling the gap between the upper and lower bounds is left open for future research.
@@ 2.1 Lattice rules
$P_{N,\boldsymbol{z}}:=\left\{\boldsymbol{x}_{n}=\left(\left\{\frac{nz_{1}}{N}\right\},\ldots,\left\{\frac{nz_{d}}{N}\right\}\right)\,\mid\,0\leq n<N\right\}.$
@@ 2.3 Lattice algorithm for approximation
As a reference value, we use the initial error where we approximate the function with 0, i.e.,
@@
see [8, Section 3.4] and [9, Remark 13.4], where $\zeta(x):=\sum_{i=1}^{\infty}i^{-x}$ for $x>1$ denotes the Riemann zeta function. When $\alpha$ is an integer, we further have
@@ 3 Construction algorithm
In the step 3.(b), we need to arrange the integers $1,\ldots,N-1$ such that the corresponding value $R_{N,s,\alpha,\boldsymbol{\gamma}}$ is listed in ascending order and then pick one of the first $\lceil\tau(N-1)\rceil$ integers. This arrangement may not be unique if some of the integers yield the same value of $R_{N,s,\alpha,\boldsymbol{\gamma}}$. However, we can always make the ordering unique by further arranging those integers in ascending order. Moreover, as shown in Remark 2.5, when $\alpha$ is an integer, $R_{N,s,\alpha,\boldsymbol{\gamma}}$ has a computable formula as
@@
For the induction step, let $\boldsymbol{z}_{s-1}$ be the (s1)(s-1)-dimensional vector drawn by the first s1s-1 steps of the algorithm. For any $z_{s}\in\{1,\ldots,N-1\}$, by separating the cases with $\ell_{s}=0$ and $\ell_{s}\neq 0$, we have
@@
$\quad\quad\quad\times\sum_{\boldsymbol{h}\in\mathbb{Z}^{s-1}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{s-1}\\ \boldsymbol{\ell}\cdot\boldsymbol{z}_{s-1}\equiv-\ell_{s}z_{s}\ (\mathrm{mod}\ N)\end{subarray}}\frac{1}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}$
@@
$\quad\quad\quad\times\sum_{\boldsymbol{h}\in\mathbb{Z}^{s-1}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{s-1}\\ \boldsymbol{\ell}\cdot\boldsymbol{z}_{s-1}\equiv-\ell_{s}z_{s}\ (\mathrm{mod}\ N)\end{subarray}}\frac{1}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}.$
@@
$\quad\quad\quad\times\sum_{\boldsymbol{h}\in\mathbb{Z}^{s-1}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{s-1}\\ \boldsymbol{\ell}\cdot\boldsymbol{z}_{s-1}\equiv-\ell_{s}z_{s}\ (\mathrm{mod}\ N)\end{subarray}}\frac{1}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}))^{1/\lambda}}$
@@
$\quad\quad\quad\times\sum_{\boldsymbol{h}\in\mathbb{Z}^{s-1}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{s-1}\\ \boldsymbol{\ell}\cdot\boldsymbol{z}_{s-1}\equiv 0\ (\mathrm{mod}\ N)\end{subarray}}\frac{1}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}))^{1/\lambda}}$
@@
$\quad\quad\quad\times\sum_{\boldsymbol{h}\in\mathbb{Z}^{s-1}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{s-1}\\ \boldsymbol{\ell}\cdot\boldsymbol{z}_{s-1}\not\equiv 0\ (\mathrm{mod}\ N)\end{subarray}}\frac{1}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}))^{1/\lambda}}$
@@ 4.1 Our algorithm
where $N,\boldsymbol{z},\boldsymbol{\Delta}$ are the random variables and $\boldsymbol{\Delta}$ follows the uniform distribution over $0,1)^{d}$, and
@@
[…]
@@
Although our randomized method is based on Algorithm 3.1, it can be substituted with Algorithm 2.4 in [7]. The difference lies in the error criterion being minimized. Although we omit a detailed discussion in this paper, the randomized method based on Algorithm 2.4 in [7] leads to an upper bound of the same order as the one we prove in Theorem 4.5 and Corollary 4.6; namely, the $M^{-\alpha(2\alpha+1)/(4\alpha+1)+\varepsilon}$ decay of the worst-case RMSE. The major advantage of using Algorithm 3.1 lies in its superior performance in the deterministic setting: for each realization of the number of points $N$ and the generating vector $\boldsymbol{z}$, the corresponding deterministic, single rank-1 lattice-based approximation algorithm yields a better convergence rate for the worst-case $L_{2}$-approximation error. The difference in convergence rates between the two deterministic lattice-based methods is discussed in detail in [9, Chapter 13], particularly in the paragraphs titled “A direct approach” and “Improving the convergence rate.” Consequently, our randomized method based on Algorithm 3.1 is expected to exhibit more stable and reliable approximation behavior.
@@ 4.2 Error Analysis
[…]
@@
$H_{M}:=\inf_{1/2\leq\lambda<\alpha}\left(\frac{2}{(1-\tau)M}\prod_{j=1}^{d}\left(1+2^{2\alpha+2}\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{2}\right)^{\lambda}.$ (8)
@@
$M\geq\inf_{1/2\leq\lambda<\alpha}\frac{2}{1-\tau}\prod_{j=1}^{d}\left(1+2^{2\alpha+2}\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{2}$ (9)
@@
    There exists a constant $c>0$ such that, for all $\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}$ we have
@@
$\leq\frac{1}{T}+C_{\alpha,\beta,\lambda,\tau}\frac{T^{1+1/(2\lambda)}}{M^{2\lambda-\lambda\beta/\alpha+1}}\prod_{j=1}^{d}\left(1+2^{2\alpha+2}\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{4\lambda-2\lambda\beta/\alpha}\max(1,2^{\beta}\gamma_{j}^{\beta/\alpha}),$
for any $\lambda\in[1/2,\alpha)$ and $\beta\in(0,1]$, with a constant $C_{\alpha,\beta,\lambda,\tau}>0$ independent of $M,d,\boldsymbol{\gamma},T$.
@@
$\left(e^{\mathrm{rms}\text{-}L_{2}\text{-}\mathrm{app}}_{d,\alpha,\boldsymbol{\gamma}}(A^{\mathrm{rand}}_{N,\boldsymbol{z},\boldsymbol{\Delta},\mathcal{A}_{d}(T)})\right)^{2}\leq\frac{1}{T}+\sum_{\boldsymbol{h}\in\mathcal{A}_{d}(T)}\sup_{\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}\frac{\omega(\boldsymbol{\ell})}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}.$
@@
Using this and the first assertion of Lemma 4.4, together with the elementary inequality $\log(1+x)\leq x^{\beta}/\beta$ for $x>0$ and $0<\beta\leq 1$ with $\log$ denoting the natural logarithm, we have
$B_{M}^{2}$ $=\sum_{\boldsymbol{h}\in\mathcal{A}_{d}(T)}\sup_{\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}\frac{\omega(\boldsymbol{\ell})}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}$
$\leq\frac{c}{\tau M}\sum_{\boldsymbol{h}\in\mathcal{A}_{d}(T)}\sup_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\geq H^{-1}_{M}\end{subarray}}\frac{\log\left(1+\|\boldsymbol{\ell}\|_{\infty}\right)}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}$
$\leq\frac{c}{\beta\tau M}\sum_{\boldsymbol{h}\in\mathcal{A}_{d}(T)}\sup_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\geq H^{-1}_{M}\end{subarray}}\frac{\|\boldsymbol{\ell}\|_{\infty}^{\beta}}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}$
$\leq\frac{c}{\beta\tau M}\sum_{\boldsymbol{h}\in\mathcal{A}_{d}(T)}\sup_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\geq H^{-1}_{M}\end{subarray}}\frac{\left(r_{\alpha,\boldsymbol{\gamma}}^{2}(\boldsymbol{\ell})\right)^{\beta/(2\alpha)}}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})},$
@@
$B_{M}^{2}$ $\leq\frac{c}{\beta\tau M}\prod_{j=1}^{d}\max(1,2^{\beta}\gamma_{j}^{\beta/\alpha})\sum_{\boldsymbol{h}\in\mathcal{A}_{d}(T)}\sup_{\begin{subarray}{c}\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\geq H^{-1}_{M}\end{subarray}}\frac{\left(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})\right)^{\beta/\alpha}}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}))^{2-\beta/\alpha}}$